OpenStudy (anonymous):

I am really confused. Solve the following system of linear equations by substitution: -2x+6y=24 3x-2y=-1

OpenStudy (anonymous):

Divide 1st equation by -2. x-3y=-12 x=3y-12 substitute it into the 2nd eq-n 3(3y-12)-2y=-1 9y-36-2y=-1 7y-36=-1 7y=-1+36=35 y=5

sam (.sam.):

-2x+6y=24 3x-2y=-1 ------------------- y=(24+2x)/6 3x-2(24+2x/6)=-1 -------------------- x=3 y=5

OpenStudy (anonymous):

use y value in 1st eq-n x-3*5=-12 x-15=-12 x=-12+15=3 (3,5)